# A Novel Approach to Industrial Defect Generation through Blended Latent Diffusion Model with Online Adaptation

### v1

@@ Abstract
Effectively addressing the challenge of industrial Anomaly Detection (AD) necessitates an ample supply of defective samples, a constraint often hindered by their scarcity in industrial contexts. This paper introduces a novel algorithm designed to augment defective samples, thereby enhancing AD performance. The proposed method tailors the blended latent diffusion model for defect sample generation, employing a diffusion model to generate defective samples in the latent space. A feature editing process, controlled by a ”trimap” mask and text prompts, refines the generated samples. The image generation inference process is structured into three stages: a free diffusion stage, an editing diffusion stage, and an online decoder adaptation stage. This sophisticated inference strategy yields high-quality synthetic defective samples with diverse pattern variations, leading to significantly improved AD accuracies based on the augmented training set. Specifically, on the widely recognized MVTec AD dataset, the proposed method elevates the state-of-the-art (SOTA) performance of AD with augmented data by 1.5%, 1.9%, and 3.1% for AD metrics AP, IAP, and IAP90, respectively. The implementation code of this work can be found at the GitHub repository https://github.com/GrandpaXun242/AdaBLDM.git
@@ I Introduction
In a practical manufacturing workflow, obtaining defective samples is considerably more challenging compared to acquiring defect-free ones. Consequently, the majority of recently proposed algorithms for industrial defect inspection, such as [1, 2, 3, 4, 5, 6, 7], address the issue as an Anomaly Detection (AD) problem—an established machine learning challenge [8, 9].
In these industrial AD algorithms, the assumption is made that all anomaly-free samples, be they complete images or image patches, belong to a single distribution. Meanwhile, defective samples are identified as “outliers”. The strategic choice of adopting an Anomaly Detection approach for industrial defect inspection is evident: the absence of the need for defective samples during the training stage makes these AD-based algorithms inherently compatible with practical manufacturing scenarios.
@@
On the flip side, achieving this compatibility comes at the expense of an extremely imbalanced data distribution, a characteristic that is deemed unacceptable in the context of most discriminative algorithms. Consequently, some researchers suggest the generation of artificial defect patterns using conventional approaches [1, 3, 5] to facilitate subsequent discriminative learning processes.
More recently, state-of-the-art (SOTA) algorithms [10, 11, 12] highlight that, in addition to synthetic defects, incorporating a few real anomalous samples during the training stage can significantly enhance Anomaly Detection (AD) accuracy. This approach holds the promise of achieving even higher performance when superior simulated defective patterns are integrated into training. Pioneering works [13, 14] have already demonstrated the benefits of generating these ”lifelike” defects using more sophisticated methods. The first two rows of Fig. 1 illustrate the existing two styles of defect generation.
In this paper, we propose the integration of the cutting-edge AI generation algorithm, namely, the Diffusion Probabilistic Model (DPM) [15], into the realm of defect generation. Specifically, we leverage the Blended Latent Diffusion Model [16, 17, 18] as our foundational framework. To enhance the generation of defective samples, we implement three key modifications to the original Blended Latent Diffusion Model (BLDM). Firstly, we design a novel “defect trimap” to delineate the target object mask and the defective regions on the generated image, incorporating it as a new form of controlling information alongside the language prompts within the stable diffusion model. Secondly, to ensure the authenticity of the generated samples, a cascade “editing” stage is introduced for the BLDM, operating in both the latent and pixel space. Finally, an innovative online adaptation of the image encoder is proposed to refine the quality of the generated images. Consequently, the highly customized BLDM algorithm, referred to as AdaBLDM in this paper, exhibits the ability to generate defective samples with significantly higher image quality and pattern variations compared to those produced by state-of-the-art (SOTA) methods. The third row of Fig. 1 provides a brief illustration of the proposed BLDM-based method. In summary, the contributions of this paper are threefold, as outlined below.
@@ II-A Synthetic Defect Generation
As introduced in Sec. I, due to a scarcity of defective samples, many researchers [1, 2, 3, 5, 6] are “compelled” to generate synthetic defective samples for better AD performance. The most straightforward way to mimic defects is to “paste” anomaly pixels on the normal images as a special case of data augmentation. Baseline methods like [20], CutPaste [21], and [1] randomly cut regions from normal images and paste them to the “incorrect” places as artificial defects. Crop&Paste [22] and PRN[23] crop defect areas from genuine defective images and paste them onto defect-free images. Those algorithms, though achieve better AD performance compared with the standard one-class AD approaches, can not create new defect patterns and thus likely lead to overfitting problems. To increase the variations of anomalies, DRAEM [3], DeSTseg [6], MegSeg [5] and ReSynthDetect [24] employ extra datasets combined with Berlin noise to fabricate defects. However, in these cases, the distribution of fabricated defects differs from the real one and thus the performance gain can not be ensured.
@@
Encouraged by the great success of AIGC (AI Generated Contents) methods including GAN (Generative Adversarial Network) algorithms [25] and Diffusion Models [26], some efforts have been made to generate more natural-looking and various simulated defects. In particular, SDGAN [27] employs two generators to switch the defective and non-defective status for sample images. SDGAN can generate a high-quality, diverse dataset of steel surface with both defective and defect-free images, thus higher classification performance is obtained. Similarly, Defect-GAN [28] can convert non-defective images into defective ones to enhance the learning process of the defect classifier. However, the above two GAN-based generation algorithms can not generate well-defined defective regions. Consequently, they could boost the performance of defect classification while can hardly lift the defect segmentation accuracy which is crucial in many real-world applications.
@@
To the best of our knowledge, [13] and [14] are the only two exiting works that generate both the synthetic defective images as well as the corresponding pixel-level labels. In specific, the DCDGANc algorithm proposed in [13] design a novel GAN-based method to mimic real defects and the defects are fused with defect-free images via an improved Poisson blending algorithm. On the other hand, [14] introduced DFMGAN, stems from StyleGANv2 [29], to generate defect images and defect masks by using the proposed defect-aware residual blocks. Although the above two GAN-based methods can generate high-quality images with defects, they fall short in generating the defect mask strictly aligned to the generated defective pixels on the image. The misalignment stems from the disparate nature of generative methods which is a blessing in image diversity whereas it could significantly mislead the defect segmentation model in the scenario of AD. In contrast, in this paper, we propose a novel defect generation method based on Diffusion Models. We tame this AIGC algorithm to generate diverse, realistic but accurately-controlled defective image samples via a series of novel modifications.
@@ II-B Diffusion Probabilistic Models for Image Editing
Compared with the more traditional generation methods like Generative Adversarial Networks (GANs) [30], Variational AutoEncoders (VAEs) [31] and flow-based algorithms [32], Diffusion Probabilistic Models (DPMs) [15, 33, 34, 35] enjoy a higher quality of the synthetic images and impressive stability of training. Accordingly, during the past few years, DPM-based approaches, such as Stable Diffusion (Stability AI) [36], DALL-E2 (OpenAI) [37], and Imagen (Google) [38], have achieved state-of-the-art performances in the task of content generation.
In the subdomain of image editing, DPM-based methods also play the most important role nowadays. In particular, the Latent Diffusion Model [16] conducts the image generation or editing in the latent space with a much less feature dimensionality and thus achieve much higher editing efficiency. RePaint [39] fully leverages a pre-trained DDPM [34] to perform image editing by sampling given pixels during the reverse diffusion stage. DCFace [40] employs two feature encoders to control the Diffusion Model, facilitating facial image editing. [41] proposes the ControlNet to govern existing Diffusion models to achieve a better-controlled synthesis. [17, 18] enables image editing by making modifications in either the pixel space or feature space. [42] introduces sparse text optimization to achieve better image editing.
Those methods, though achieve SOTA performance in general image editing tasks, the fuzzily governed synthesis easily fails to generate the image regions aligned with pixel-level accuracy. This can potentially decrease segmentation model performance. In this paper, we highly customize the Blended Latent Diffusion Model (BLDM) [18] algorithm for generating the precisely located defect regions for industrial defect detection and localization.
Note that this work employs the diffusion model for generating synthetic datasets, just like the pioneering work [15, 43] do for more general purposes. Our method is orthogonal to those algorithms that use DPM to predict anomaly regions [44, 45, 46, 47].
@@ III-A Task Definition and Preliminaries
Following the conventional setting of anomaly detection, an anomaly-free (OK) sample set $\mathcal{X}_{OK}=\{\mathbf{x}^{i}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{OK}\}$ is available for training. Furthermore, as introduced in Sec. I, a few defective (NG) samples $\mathcal{X}_{NG}=\{\mathbf{x}^{i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{NG}\}$ and the corresponding anomaly masks $\mathcal{M}_{NG}=\{\mathbf{M}^{i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N_{NG}\}$ are given as the “seeds” of defect generation and usually $N_{NG}\ll N_{OK}$. The task of this paper is to generate the synthetic training set of defective samples as
@@
where $\mathcal{X}^{\ast}_{NG}=\{\mathbf{x}^{\ast i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ and $\mathcal{M}^{\ast}_{NG}=\{\mathbf{M}^{\ast i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ comprise the obtained synthetic sample images and the corresponding label masks, respectively. Furthermore, the defect generation algorithm usually sets $N^{\ast}_{NG}\gg N_{NG}$ to ensure higher sample variations.
@@
where $\boldsymbol{\mu}_{\theta}(\cdot)$ and $\boldsymbol{\Sigma}_{\theta}(\cdot)$ are the two unknown functions that are usually estimated as deep models. In the test phase, given different random Gaussian noise samples $\mathbf{x}^{i}_{T},i=1,\dots,N_{tst}$, one could obtain high quality and various samples $\mathbf{x}^{i}_{0},\forall i$ (usually RGB images) from the same domain of the training samples. Akin to the above diffusion process, we could also generate lifelike anomalous images in $\mathcal{X}^{\ast}_{NG}$ based on a diffusion model which has been well learned on the samples sets $\mathcal{X}_{OK}$ and $\mathcal{X}_{NG}$.
@@ III-B1 Latent diffusion model with controlling information
In this paper, we employ the Latent Diffusion Model (LDM) as the backbone of our method encouraged by its praiseworthy advantages [16]. As the name indicating, the forward diffusion and reverse diffusion processes of LDM are all performed in the latent space. In general, given the input image $\mathbf{x}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, its corresponding latent feature $\mathbf{z}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained using the encoder function $\Omega(\cdot)$ as $\mathbf{z}=\Omega(\mathbf{x})$ while the image can also be recovered via the decoder function $\Phi(\cdot)$ as $\mathbf{x}=\Phi(\mathbf{z})$. Based on the latent feature $\mathbf{z}$ and some simplifications  [^1] Readers are referred to [16] and [33] for details from Eq. 3, the reverse conditional probability in LDM is defined as
@@ III-B2 Linguistic prompts
First, a linguistic prompt is obtained via changing the keywords of a following template sentence
@@
where $obj$ and $def$ denote to the category of the current object (as the capsule in Fig. 2) and defect type (as the black contamination on the capsule), respectively. One can see this template design is straightforward and thus the prompt is always easy to obtain. In this work, the sentence $y$ is then processed by the parameter-frozen text encoder $\tau(\cdot):\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{lang}}$ which is proposed in the CLIP algorithm [48]. In particular, as shown in Fig. 2, the linguistic controlling information is fed into the generation model $\epsilon_{\theta}(\cdot)$ and the generation process is influenced by the K-Q-V attention operations [16].
III-B3 Defect trimaps
Additionally, recall that in practical anomaly detection tasks, the defect-free samples are usually sufficient in terms of both amount and variation, we thus propose to generate defective samples under the guidance of the randomly selected defect-free samples. In a nutshell, a “trimap” $\Gamma\in\mathbb{R}^{H_{x}\times W_{x}}$ is designed to specify the desired location of the generated object and the defect, as shown in Fig. 2. To obtain a proper trimap, we firstly estimate the foreground region i.e., the interested object, by using the method proposed in [11] and store the foreground mask into $\mathbf{F}\in\mathbb{B}^{H_{x}\times W_{x}}$ with $\mathbf{F}(x,y)=1$ indicating a foreground at the coordinate $[x,y]$. Secondly, a synthetic defect mask is generated based on the genuine anomaly mask set $\mathcal{M}_{NG}$ as

### v2

@@ Abstract
Effectively addressing the challenge of industrial Anomaly Detection (AD) necessitates an ample supply of defective samples, a constraint often hindered by their scarcity in industrial contexts. This paper introduces a novel algorithm designed to augment defective samples, thereby enhancing AD performance. The proposed method tailors the blended latent diffusion model for defect sample generation, employing a diffusion model to generate defective samples in the latent space. A feature editing process, controlled by a “trimap” mask and text prompts, refines the generated samples. The image generation inference process is structured into three stages: a free diffusion stage, an editing diffusion stage, and an online decoder adaptation stage. This sophisticated inference strategy yields high-quality synthetic defective samples with diverse pattern variations, leading to significantly improved AD accuracies based on the augmented training set. Specifically, on the widely recognized MVTec AD dataset, the proposed method elevates the state-of-the-art (SOTA) performance of AD with augmented data by 1.5%, 1.9%, and 3.1% for AD metrics AP, IAP, and IAP90, respectively. The implementation code of this work can be found at the GitHub repository https://github.com/GrandpaXun242/AdaBLDM.git
@@ I Introduction
In a practical manufacturing workflow, obtaining defective samples is considerably more challenging compared to acquiring defect-free ones. Consequently, the majority of recently proposed algorithms for industrial defect inspection, such as [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], address the issue as an Anomaly Detection (AD) problem—an established machine learning challenge [11, 12]. Among these industrial AD algorithms, the assumption is made that all anomaly-free samples, be they complete images or image patches, belong to a single distribution. Meanwhile, defective samples are identified as “outliers”. The strategic choice of adopting an Anomaly Detection approach for industrial defect inspection is evident: the absence of the need for defective samples during the training stage makes these AD-based algorithms inherently compatible with practical manufacturing scenarios.
@@
Nonetheless, achieving this compatibility comes at the expense of an extremely imbalanced data distribution, a characteristic that is deemed unacceptable in the context of most discriminative algorithms. Consequently, some researchers suggest the generation of artificial defect patterns using conventional approaches [1, 3, 5] to facilitate subsequent discriminative learning processes. Recently, state-of-the-art (SOTA) algorithms [13, 14, 15] suggest that incorporating a few real anomalous samples during the training stage can significantly enhance the AD accuracy if these samples are combined with synthetic samples. For example, some pioneering works [16, 17] demonstrated the benefits of generating these ”lifelike” defects using more sophisticated methods. The first two rows of Fig. 1 illustrate two different types of defect generation.
In this paper, we propose the integration of the cutting-edge AI generation algorithm, namely the Diffusion Probabilistic Model (DPM) [18], into the realm of defect generation. Specifically, we leverage the Blended Latent Diffusion Model [19, 20, 21] to create our framework. To enhance the generation of defective samples, we innovate the Blended Latent Diffusion Model (BLDM)[21] with three modules. First, we design a novel “defect trimap” to delineate the target object mask and the defective regions on the generated image, incorporating it as a new form of controlling information alongside the language prompts within the stable diffusion model. Second, to ensure the authenticity of the generated samples, a cascaded “editing” stage is introduced into the BLDM, at both the latent and pixel space. Finally, an innovative online adaptation of the image encoder is proposed to refine the quality of the generated images. Consequently, the customized BLDM algorithm, referred to as AdaBLDM in this paper, is capable of generating higher-quality defective samples with pattern variations. The third row of Fig. 1 provides a brief illustration of the proposed BLDM-based method. In summary, the contributions of this paper are threefold, as outlined below.
@@ II-A Synthetic Defect Generation
Due to the scarcity of defective samples, generating synthetic defective samples becomes a rule of thumb among researchers to achieve satisfactory performance [1, 2, 3, 5, 6, 23]. A straightforward way akin to data augmentation is mimicking the defects by “pasting” anomalous pixels on the normal images. For example, methods in the regime [24, 25, 1] randomly cut regions from normal images and paste them to the “incorrect” places as artificial defects. Crop&Paste [26] and PRN[27] crop defect areas from genuine defective images and paste them onto defect-free images. Those algorithms, though achieve better AD performance compared with the standard one-class AD approaches, can not create new defect patterns and thus likely lead to overfitting problems. To increase the variations of anomalies, DRAEM [3], DeSTseg [6], MegSeg [5] and ReSynthDetect [28] employ extra datasets combined with Berlin noise to fabricate defects. However, in these cases, the distribution of fabricated defects differs from the real one and thus the performance gain can not be ensured.
@@
Inspired by the potent AIGC (AI Generated Contents) methods such as Generative Adversarial Network (GAN) based algorithms [29, 30, 31] and diffusion model-based algorithms [32], some attempts have been devoted to generating more natural-looking simulated defects. In particular, SDGAN [30] employs two generators to switch the defective and non-defective status for sample images. As a result, SDGAN [30] can generate a high-quality, diverse dataset of steel surfaces with both defective and defect-free images. In a similar vein to SDGAN [30], Defect-GAN [33] converts non-defective images into defective ones to enhance the learning process of the defect classifier. However, the above two GAN-based generation algorithms cannot generate defective regions that are aligned precisely as required. Despite that they could enhance the performance of defect classification, they cannot lift the defect segmentation accuracy which is crucial in many real-world applications.
@@
More recently, [16] and [17] proposed to generate both synthetic defective images and the corresponding pixel-level labels. In specific, the DCDGANc algorithm [16] designed a novel GAN-based method to mimic real defects, which can be fused with defect-free images via an improved Poisson blending algorithm. At the same time, [17] introduced DFMGAN, stemming from StyleGANv2 [34], to generate defective images and defective masks by using the proposed defect-aware residual blocks. Although the above two GAN-based methods can generate high-quality images with defects, they fall short in generating the defective mass strictly aligned to the generated defective pixels on the image. Such misalignment is ascribed to the disparity between the image-specific generative methods and the challenging scenario of AD, and thus significantly misleads the defective segmentation model in AD. In contrast, in this paper, we propose a novel defect generation method based on Diffusion Models. We tame this general-purpose algorithm to generate diverse, realistic but accurately-controlled defective image samples via a series of novel modifications.
@@ II-B Diffusion Probabilistic Models for Image Editing
In the last few years, researchers have proposed various approaches to realize high-quality image generation. The off-the-shelf tools include Generative Adversarial Networks (GANs) [35], Variational AutoEncoders (VAEs) [36], the flow-based algorithms [37], and Diffusion Probabilistic Models (DPMs) [18, 38, 39, 40]. In specific, DPMs have demonstrated higher qualities of the synthetic images and impressive stability of training. Accordingly, the DPM-based approaches, such as Stable Diffusion (Stability AI) [19], DALL-E2 (OpenAI) [41], and Imagen (Google) [42], have achieved state-of-the-art performances in the task of content generation.
DPM-based methods also play an important role in the remit of image editing. In particular, the Latent Diffusion Model [19] conducts the image generation or editing in the lower-dimensional latent space and thus can achieve higher efficiency; RePaint [43] fully leverages a pre-trained DDPM [39] to perform image editing by sampling given pixels during the reverse diffusion stage; DCFace [44] employs two feature encoders to control the Diffusion Model to facilitate facial image editing; [45] proposes the “ControlNet” to govern existing Diffusion models to achieve a better-controlled synthesis; [20, 21] enable image editing by making modifications in either the pixel space or feature space; [46] propose the ”Dual-Cycle Diffusion” to generate an unbiased mask to guide image editing. While these methods demonstrate state-of-the-art (SOTA) performances in general image editing tasks, they can hardly produce image regions that are precisely aligned at the pixel level due to the inherently fuzzy nature of the reverse diffusion process. This nuisance can potentially decrease the segmentation capacity in anomaly detection.
In this paper, we customize the cutting-edge Blended Latent Diffusion Model (BLDM) [21] to generate precisely aligned synthetic regions for industrial defect detection and localization. It is important to note that this work employs the diffusion model for generating synthetic images or image regions, just similar to the pioneering work [18, 47] proposed for the general editing tasks. Our method is irrelevant to the algorithms that use DPM to directly predict anomaly regions on the test images [48, 49, 50, 51].
@@ III-A Task Definition and Preliminaries
Following the conventional setting of anomaly detection, an anomaly-free (OK) sample set $\mathcal{X}_{OK}=\{\mathbf{x}^{i}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{OK}\}$ is available for training. Furthermore, as introduced in Sec. I, a few defective (NG) samples $\mathcal{X}_{NG}=\{\mathbf{x}^{i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{NG}\}$ and the corresponding anomaly masks $\mathcal{M}_{NG}=\{\mathbf{M}^{i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N_{NG}\}$ are given as the “seeds” of defect generation and usually $N_{NG}\ll N_{OK}$. The goal of the proposed method is to generate a number of training samples with synthetic defective regions, i.e.,
@@
where $\mathcal{X}^{\ast}_{NG}=\{\mathbf{x}^{\ast i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ and $\mathcal{M}^{\ast}_{NG}=\{\mathbf{M}^{\ast i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ comprise the obtained synthetic samples and the corresponding defect masks, respectively. Furthermore, usually we set $N^{\ast}_{NG}\gg N_{NG}$ to ensure high pattern variation of the generated samples.
@@
where $\boldsymbol{\mu}_{\theta}(\cdot)$ and $\boldsymbol{\Sigma}_{\theta}(\cdot)$ are the two unknown functions that are usually approximated as deep models. In the test phase, given different random Gaussian noise samples $\mathbf{x}^{i}_{T},i=1,\dots,N_{tst}$, one could obtain high quality and various samples $\mathbf{x}^{i}_{0},\forall i$ (usually RGB images) from the same domain of the training samples. Akin to the above diffusion process, we could also generate lifelike anomalous images in $\mathcal{X}^{\ast}_{NG}$ based on a diffusion model which has been well learned on the samples sets $\mathcal{X}_{OK}$ and $\mathcal{X}_{NG}$.
@@ III-B1 Latent diffusion model with controlling information
In this paper, we employ the Latent Diffusion Model (LDM) as the backbone of our method encouraged by its praiseworthy advantages [19]. As the name indicating, the forward and reverse diffusion processes of LDM are both performed in the lower-dimentional latent space. Given an input image $\mathbf{x}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, its corresponding latent feature $\mathbf{z}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained using the encoder function $\Omega(\cdot)$ as $\mathbf{z}=\Omega(\mathbf{x})$ while the image can also be recovered from $\mathbf{z}$ via the decoding operation as $\mathbf{x}=\Phi(\mathbf{z})$. Based on the latent feature $\mathbf{z}$ and some simplifications  [^1] Readers are referred to [19] and [38] for more details from Eq. 3, the reverse conditional probability in latent space is revised as
@@ III-B2 Linguistic prompts
As an effective module, linguistic prompts are introduced to control and stabilize the generation process of diffusion models [19, 21]. In this work, the linguistic prompt $y$ is obtained by specifying the keywords in the following template sentence
@@
where the keywords $obj$ and $def$ denote the category of the current object (as the capsule in Fig. 2) and defect type (as the black contamination on the capsule), respectively. One can see this template design is straightforward and thus the prompt is always easy to obtain. The sentence $y$ is then processed by the parameter-frozen text encoder $\tau(\cdot):\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{lang}}$ of the CLIP model [52]. In particular, as shown in Fig. 2, the linguistic controlling information is fed into the generation model $\epsilon_{\theta}(\cdot)$ and the generation process is governed via a K-Q-V attention mechanism proposed in [19].
III-B3 Defect trimap prompts
Recall that in practical anomaly detection tasks, the defect-free samples are usually sufficient in terms of both amount and pattern variation, we thus propose to generate defective samples based on defect-free prototypes, rather than generating them from scratch. In a nutshell, a “trimap” $\Gamma\in\mathbb{R}^{H_{x}\times W_{x}}$ is designed to specify the desired location of the generated object and the defect, as shown in Fig. 2. To obtain a proper trimap, we firstly estimate the foreground region  [^2] The area corresponding to the interested object on a randomly selected defect-free image $\mathbf{x}_{OK}$ by using the method proposed in [14] and save the foreground mask as $\mathbf{F}\in\mathbb{B}^{H_{x}\times W_{x}}$ with $\mathbf{F}(x,y)=1$ indicating a foreground pixel at the coordinate $[x,y]$. Secondly, a synthetic defect mask is created based on a genuine anomaly mask randomly selected from set $\mathcal{M}_{NG}$. Formally, the generation of the novel anomaly mask writes: